Example 2 - Part 1 The annual salary for a certain job has a normal distribution with a mean of $54,000 and a standard deviatCurve 3 la 20 Curve 4 lo 20 Curve 5 -30 -la la d. P(x = 54, 215) = P(= = = = Thorn in min) ban monnaio of random 054 915
Answer #1

solution:- M = 54000 o sooo 2) a lix= M- 54000 b) Tf - SG = 5ooo 50o loo Z = X-1 - 0.43 54 215 - 54000 500 ur M uto d) plo 54